OpenStudy (anonymous):

What polynomial has roots of -6, -4, and 1 ? x3 - 9x2 - 22x + 24 x3 - x2 - 26x - 24 x3 + x2 - 26x + 24 x3 + 9x2 + 14x - 24

OpenStudy (solomonzelman):

Have you ever done this type of problems before? If what, what did you do ?

OpenStudy (cwrw238):

(x+6)(x+4)(x-1) = 0

OpenStudy (anonymous):

ok then what do i do? @cwrw238

OpenStudy (anonymous):

Would I foil @cwrw238

OpenStudy (cwrw238):

you can foil (x+4)(x-1) if you like then multiply this by (x + 6)

OpenStudy (cwrw238):

this will give you (x + 6)(x^2 + 3x - 4) now use the distributive law multiply the second bracket by x then by + 6 and simplify to give one of the above options

OpenStudy (campbell_st):

well an easier method may be to use the sum and product of the roots of a cubic \[ax^3 + bx^2 + cx + d = 0\] then using the connection between roots and coefficients sum of the roots \[-\frac{b}{a} = -6 + (-4) + 1 \] in this question a = 1 so -b = -9 hence b = 9 then the sum of the porduct of pairs \[\frac{c}{a} = -6 \times -4 + -6 \times 1 + -4 \times 1 = 14\] then c = 14 and lastly the product of the roots \[\frac{-d}{a} = -4 \times -6 \times 1\] -d = 24 so d = -24 so the substitute a = 1, b = -9 c = 14 and d = -24 into the general form at the top

OpenStudy (anonymous):

General method: substitute these values: -6, -4, 1 into these expressions to see which ones equal zero. Only the third expression has (1) as root. Only the second expression has (-4) as root Only the forth expression has (-6) as root.

OpenStudy (campbell_st):

another option is to use the factor theorem find f(-4) f(-6) f(1) ther equation where all 3 are zero is your answer... which is D